Solve the following equations and show work please?
1.) √x-1 = 3
2.) √x^3 = 8
3.) 3√x^2 = 4

Answer by mukhopadhyay(490)

1.) √x-1 = 3
=> √x = 4
=> x = 16 (squaring both sides)
..........
2.) √x^3 = 8
=> x^3 = 64 (squaring both sides)
=> x = 4;
......
3.) 3√x^2 = 4
=> x^(2/3) = 4
=> [x^(2/3)]^3 = 4^3
=> x^2 = 64
=> x = 8 (negative ignored because 3√x^2 is a positive number)